OpenStudy (superflytyguy):

I really need help (will give medal) A sound wave is modeled with the equation y = 1/4 cos 2 pi/3 theta a. Find the period. Explain your method. b. Find the amplitude. Explain your method. c. What is the equation of the midline? What does it represent? I really need help on this one!

OpenStudy (superflytyguy):

\[y = \frac{ 1 }{ 4 } \cos \frac{ 2\pi }{ 3 } \Theta\]

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

period \[\cos(bx)\] is \[\huge \frac{2\pi}{b}\] in your case \(\huge b=\frac{2\pi}{3}\)

OpenStudy (misty1212):

divide and you will get it

OpenStudy (misty1212):

b) amplitude is the number out front i.e \(\frac{1}{4}\)

OpenStudy (misty1212):

c) since you are not adding or subtracting anything outside, the "midline" is the x axis \(y=0\)

OpenStudy (superflytyguy):

2pi / 3 = 2.094 or about 2.1

OpenStudy (misty1212):

lol no

OpenStudy (misty1212):

the period is not \(b\) it is \(\frac{2\pi}{b}\) and the question is no asking for you to convert to decimal form

OpenStudy (misty1212):

\[\huge \frac{2\pi}{\frac{3\pi}{2}}\] is what you need to compute put the calculator away it will not help you here

OpenStudy (superflytyguy):

I see, sorry I'm not the best at math. English is my strong suit.

OpenStudy (misty1212):

you know how to compute that? "invert and multiply" as the math teachers say

OpenStudy (superflytyguy):

Thank you for your help!